a cellular phone company charges a $49.99 monthly fee for 600 free minutes.Each additional minute costs $.35. This month you used 750 minutes. How much do you owe?

Answer by checkley71(8403)

COST=49.99+(750-600).35
COST=49.99+150*.35 (BASE COST + 150 MINUTES @ .35 PER MINUTE)
COST=49.99+52.50
COST=$102.49 TOTAL MONTHLY COST FOR 750 MINUTES OF USE.